# A Model Predictive Capture Point Control Framework for Robust Humanoid Balancing via Ankle, Hip, and Stepping Strategies

### v2

@@ IV-C Variable Weighting Parameter for CAM control
Figure 3: Decision process of variable weighting parameter in CP–MPC.
@@
Finally, the desired centroidal moment $\bm{\tau}^{des}=[\tau_{y}^{des}\,-\tau_{x}^{des}]$ is integrated into the desired CAM at each time step $\mathbf{h}^{des}=[h_{y}^{des}\,-h_{x}^{des}]$ as follows:
@@
The desired CAM is controlled by our previously proposed HQP-CAM controller [23], which prioritizes CAM control first and performs the initial pose return task with secondary priority, resulting in improved CAM control performance. The effectiveness of the variable weighting approach will be validated in Section VI-E.
@@ V-B Parameter Decision based on CP–MPC
where $T_{k}$ represents the time step at which SSP ends within the MPC horizon, $\mathbf{z}_{i}=[z_{x,i}\,z_{y,i}]$ and $\bm{\tau}_{i}=[\tau_{y,i}\,-\tau_{x,i}]$ denote the ZMP and centroidal moment control input at each time step $i$, respectively. The derivation for $\mathbf{P}_{ssp}$ is provided in Appendix VIII.
@@ V-D Characteristic of Our Parameter Selection Approach compared to Previous Methods
In Section VI-F, the effectiveness of the parameter selection approach is analyzed through a comparison with the previous methods.
@@
Figure 5: Simulation results are presented for the robot’s response to external forces applied along both the negative x- and y-directions, shown in (a) and (b) respectively. The second and third rows depict the output of the CP–MPC, while the fourth and fifth rows show the reference foot trajectories and step time by the stepping controller.
VI Results of Simulations and Experiments
VI-A System Overview
@@
In this section, a comprehensive system overview for both the simulations and real robot experiments is presented. The humanoid robot, TOCABI, utilized in the simulations and experiments comprises a total of 33 degrees of freedom (DOFs): 16 for the arms, 12 for the legs, 3 for the waist, and 2 for the neck. Its physical dimensions are approximately 1.8 m in height, weighing around 100 kg, and with a foot size of 14 cm $\times$ 30 cm. The upper body actuators are comprised of Parker’s BLDC motors and harmonic gears, and the lower body actuators are composed of Kollmorgen’s BLDC motors and harmonic gears. The current control of the robot is performed by Elmo Motion Control’s Gold Solo Whistle servo controller, and the communication between the servo controller and the main PC is achieved via EtherCAT communication. MicroStrain’s 3DM-GX5-25 IMU is attached to the pelvis for estimating the global inclination of the robot, and ATI’s mini85 F/T sensor is mounted on each foot to measure the contact wrench of the robot. The algorithmic operation and torque command frequency of the robot are set to 2 kHz, and the walking pattern generator and CP–MPC operate at 50 Hz (shown in Fig. 2), due to the high computational load involved, through parallel threads. The walking pattern generator and CP–MPC have respective MPC time horizons of 2.5 s and 1.5 s. It should be noted that the positive direction of the x-axis represents the robot’s forward direction, while the positive direction of the z-axis is the opposite direction of the gravity vector. The simulator employed in this study is the MuJoCo simulator [49]. The qpOASES [50] was employed to solve QP optimization problems, while RBDL [51] was utilized for calculating the kinematics and dynamics of the robot.
Figure 6: Simulation results are presented for the robot’s forward walking while overcoming four unexpected objects, with snapshots and output data provided. The second and third rows represent the output of the CP–MPC, while the fourth and fifth rows show the reference foot trajectories and step time by the stepping controller.
@@
VI-B Simulations to Validate Robustness of the Proposed Method under External Forces
@@
Simulations were conducted to evaluate the balancing capability of the robot against external forces using the proposed method. The simulations involved a scenario where the robot walked in the place and external forces are applied to the robot’s pelvis link in the x- and y-directions during an SSP (left foot support). The step duration was configured to consist of an SSP of 0.6 s and a DSP of 0.3 s. Furthermore, adjustments of the footstep position and step time are not allowed 0.1 s before the end of SSP to prevent jerky motions in the swing leg.
@@
In the first simulation, an external force of 240 N was applied to the robot’s pelvis in the negative x-direction for 0.2 s. Fig. 5(a) shows snapshots and data of the robot’s response to the external force. The snapshots illustrates the overall balancing process. The robot maintains balance by employing ZMP control using the supporting legs (ankle strategy), CAM control using the upper body (hip strategy), and performing back steps (stepping strategy).
@@
The graph in Fig. 5(a) represents the outputs of the CP–MPC and stepping controller in the x-direction. At approximately 5.9 s, the robot’s CP was pushed backward, causing the CP error $\xi_{err,x}$ (green line) to increase in the negative x-direction due to the external force. To reduce the CP error, the desired ZMP $z_{x}^{des}$ (red line) is generated by the CP–MPC. Despite the attempts to reduce the CP error through the desired ZMP, it remained high due to the significant external force and the ZMP constraint of -9 cm. As a result, the CP error increased and reached up to -16.8 cm. Then, footstep adjustment $\Delta f_{x,1}$ (blue line) was generated to relax the ZMP constraint and the desired centroidal moment $\tau_{y}^{des}$ (yellow line) was also generated to further reduce the CP error. The desired ZMP was generated up to the relaxed ZMP constraint until approximately 7.4 s, while the centroidal moment was generated up to the limit of -15 Nm until approximately 7.0 s. Meanwhile, the stepping controller performed a total of two back-steps (magenta and cyan line) based on $\Delta f_{x,1}$, moving approximately -20.3 cm in conjunction with an optimal step time $T$ (orange line). As a result, the CP error is reduced to approximately zero, and the robot maintains a balance against the external force.
@@
Fig. 5(b) represents the robot’s data, including the CP–MPC and stepping controllers’ y-direction outputs. At approximately 5.9 s, the external force was applied to the robot in the negative y-direction, leading to a significant deviation in the CP error $\xi_{err,y}$. The CP–MPC attempted to decrease the increasing CP error by generating the desired ZMP up to the ZMP constraint. At approximately 6 s, footstep adjustment of -8 cm $\Delta f_{y,1}$ was generated to relax the ZMP constraint, while a desired centroidal moment $\tau_{x}^{des}$ of 15 Nm was also generated to reduce the CP error. Based on the generated $\Delta f_{y,1}$, the stepping controller adjusted the foot trajectory and reduced the step time $T$ to allow for stepping at approximately 6.1 s. Subsequently, continuous CP control using each strategy reduced the CP error from approximately -30.2 cm initially, gradually decreasing until the robot was able to maintain balance against external forces.
VI-C Simulation to Validate Robustness of the Proposed Method on Uneven Terrain
In this simulation, the robot walks over objects placed on the ground to validate the balancing performance of the proposed method on uneven terrain. Four objects, each with a length and width of 5 cm, were positioned in front of the robot with a distance of 30 cm between them. To make the robot step on each object alternately, two objects were placed on the left and two on the right with respect to the robot’s center. The thickness of the objects ranged from 1.0 cm to 2.5 cm with an increase of 0.5 cm. The robot walked a distance of 2.0 m with a step length of 0.1 m, and the step duration consisted of 0.6 s of SSP and 0.3 s of DSP. In the second row of Fig. 6, the right y-axis, highlighted in purple, depicts the CP error $\xi_{err}$ and footstep adjustment $\Delta\mathbf{f}_{1}$. This arrangement is intended to enhance the visibility of variations in CP error and footstep adjustment.
Fig. 6 shows simulation snapshots and the corresponding output data. In the snapshots, the robot was able to navigate forward while stepping on the objects in its path. The first and second objects were overcome using ZMP and stepping control (ankle and stepping strategies). For the third and fourth objects, CAM control (hip strategy) was additionally employed to navigate past them. The analysis of this process is explained based on the data presented in the graphs.
@@ VI-C Simulation to Validate Robustness of the Proposed Method on Uneven Terrain
At the start of the 9th DSP, approximately 9.9 s, the robot encountered the 1.5 cm object. To overcome the resulting disturbance, the CP–MPC generated a footstep adjustment $\Delta f_{y,1}$ of 7.3 cm at approximately 10.7 s, in contrast to the initial encounter with 1.0 cm object. By utilizing $\Delta f_{y,1}$ as a nominal value in the stepping controller, the reference swing foot position $e^{ref}_{L,R,y}$ was adjusted, enabling the robot to overcome the disturbance.
Next, at the start of the 12th DSP, approximately 12.6 s, the robot stepped on the 2 cm object. To alleviate the CP errors, the desired ZMP was generated up to the constraint in both the x- and y-directions. Additionally, $\Delta f_{y,1}$ of -9.2 cm was generated at approximately 13.1 s and $\Delta f_{x,1}$ of -12.2 cm was also generated at 13.5 s. Notably, the large CP error in the y-direction led to a centroidal moment $\tau^{des}_{x}$ of 15 Nm. The robot was able to overcome disturbances by adjusting footstep position and step time to a greater extent than in the second disturbance response, while also implementing CAM control.
Finally, at the start of the 15th DSP, approximately 15.2 s, the robot stepped on the thickest object of 2.5 cm. In contrast to the third disturbance response, balance was achieved over two steps due to a large disturbance. The desired ZMP in the x- and y-direction were generated up to each ZMP constraint in two steps. $\Delta f_{y,1}$ of 9.9 cm was generated at approximately 15.5 s, and $\Delta f_{x,1}$ was generated up to -12.4 cm at approximately 16.1 s. The centroidal moment $\tau^{des}_{x}$ in the x-direction was also generated up to the constraint. With footstep position adjustment, step time was also reduced by up to 0.48 s, allowing the robot to overcome the disturbance. In the x-direction, due to the larger range of ZMP control and footstep adjustments compared to the y-direction, only a small amount of centroidal moment $\tau^{des}_{y}$ was generated in overcoming all objects.
In conclusion, this simulation showed that appropriate strategies were executed to mitigate disturbances of various magnitudes caused by uneven terrain. The sequence in our approach, where ankle and stepping strategies are employed first, followed by the use of the hip strategy, resembles the preferred order of balance strategies in human behavior [6, 52, 53].
VI-D Simulation to Evaluate Balancing Performance based on the Different Combinations of Each Balance Strategy
Comparative simulations were conducted to evaluate how the balancing performance varies based on different combinations of balancing strategies in response to disturbances. Presented below is a summary of the implementation methods employed to combine each strategy.
  • Method 2) is composed of a combination of ZMP control and stepping control (footstep position and step time adjustment), excluding the CAM control. To achieve this, all terms related to the centroidal moment $\bm{\uptau}$ in (10) and (11) were excluded, as well as the CAM controller.
@@
VI-E Simulation to Validate the Effectiveness of the Variable Weighting Method

### v3

@@ IV-C Variable Weighting Parameter for CAM control
Figure 2: Decision process of variable weighting parameter.
@@
Finally, the desired centroidal moment $\bm{\tau}^{des}=[\tau_{y}^{des}\,-\tau_{x}^{des}]$ is computed and integrated at each time step, accumulating into the desired CAM as follows,
@@
Here, the z-axis component is set to zero, and the desired CAM $\mathbf{h}^{des}\in\mathbb{R}^{3}$ is realized through the HQP-based WBIK described in Section VI.
@@ V-B Parameter Decision based on CP–MPC
where $T_{k}$ represents the time step at which SSP ends within the MPC horizon, $\mathbf{z}_{i}=[z_{x,i}\,z_{y,i}]$ and $\bm{\tau}_{i}=[\tau_{y,i}\,-\tau_{x,i}]$ denote the ZMP and centroidal moment control input at each time step $i$, respectively. The derivation for $\mathbf{P}_{ssp}$ is provided in Appendix IX.
@@ V-D Characteristic of Our Parameter Selection Approach compared to Previous Methods
In Section VII-F, the effectiveness of the parameter selection approach is analyzed through a comparison with the previous methods.
VI HQP–based Whole-Body Inverse Kinematics
This section introduces an HQP–based WBIK to achieve the outputs from the walking pattern generator, CP–MPC, and stepping controller. The HQP-based WBIK extends our previous HQP–CAM controller [23], which was limited to upper-body CAM control, into a whole-body control framework.
The main objective of HQP–WBIK is to track the reference foot and CoM trajectories, as well as the desired CAM, while integrating initial pose recovery after CAM tracking motion into the hierarchical control framework. HQP–WBIK prioritizes CAM control over initial pose recovery by establishing strict task hierarchies to enhance balance performance.
The formulation of the HQP–WBIK is defined as follows.
$\underset{\mathbf{\dot{q}_{1}}}{\text{min}}$ $\sum_{j}{\|\mathbf{J}_{1,j}\mathbf{\dot{q}}_{1}-\mathbf{\dot{x}}^{des}_{1,j}\|^{2}_{\mathbf{w}_{\mathbf{J}}}+\|\mathbf{A}\mathbf{\dot{q}}_{1}-\mathbf{h}^{des}\|^{2}_{\mathbf{w}_{\mathbf{A}}}+\|\mathbf{\dot{q}}_{1}\|^{2}_{\mathbf{w}_{\mathbf{\dot{q}}}}}$ (21)
s. t. $\underline{\mathbf{\dot{q}}}_{1}\leq\mathbf{\dot{q}}_{1}\leq\overline{\mathbf{\dot{q}}}_{1}$ (22)
$\mathbf{K}_{q}(\underline{\mathbf{q}}_{1}-\mathbf{q}_{1})\leq\dot{\mathbf{q}}_{1}\leq\mathbf{K}_{q}(\overline{\mathbf{q}}_{1}-\mathbf{q}_{1})$
$\underset{\mathbf{\dot{q}_{2}}}{\text{min}}$ $\sum_{j}{\|\mathbf{J}_{2,j}\mathbf{\dot{q}}_{2}-\mathbf{\dot{x}}^{des}_{2,j}\|^{2}_{\mathbf{w}_{\mathbf{J}}}}+\|\mathbf{\dot{q}}_{2}\|^{2}_{\mathbf{w}_{\mathbf{\dot{q}}}}$ (23)
s. t. $\mathbf{J}_{1,j}\mathbf{\dot{q}}_{2}=\mathbf{J}_{1,j}\mathbf{\dot{q}}_{1}$ (24)
$\mathbf{A}\mathbf{\dot{q}}_{1}-\mathbf{\bm{\varepsilon}}\leq\mathbf{A}\mathbf{\dot{q}}_{2}\leq\mathbf{A}\mathbf{\dot{q}}_{1}+\mathbf{\bm{\varepsilon}}$
$\underline{\mathbf{\dot{q}}}_{2}\leq\mathbf{\dot{q}}_{2}\leq\overline{\mathbf{\dot{q}}}_{2}$
$\mathbf{K}_{q}(\underline{\mathbf{q}}_{2}-\mathbf{q}_{2})\leq\dot{\mathbf{q}}_{2}\leq\mathbf{K}_{q}(\overline{\mathbf{q}}_{2}-\mathbf{q}_{2})$
The optimization variable for the hierarchy $i\in\{1,2\}$ is the joint velocity $\mathbf{\dot{q}}_{i}\in\mathbb{R}^{n+6}$, where $n$ and $6$ represent the number of actuated joints and virtual joints, respectively. $\mathbf{J}_{i,j}$ and $\mathbf{\dot{x}}^{des}_{i,j}$ denote the Jacobian matrix and the desired velocity command for $j$-th task in the hierarchy $i$, respectively. The desired velocity command is defined as $\mathbf{\dot{x}}^{des}_{i,j}=\mathbf{\dot{x}}^{ref}_{i,j}+\mathbf{K}_{p}({\mathbf{x}}^{ref}_{i,j}-{\mathbf{x}}_{i,j})$ to minimize position or orientation error of the $j$-th task in the hierarchy $i$, where $\mathbf{K}_{p}$ is a positive feedback gains.
In the first hierarchy, the cost function (21) is composed of three cost terms. The first cost term serves as the desired velocity command tracking control, ensuring that the robot follows the reference CoM position $\mathbf{c}^{ref}\in\mathbb{R}^{3}$ and desired foot pose as $\mathbf{e}^{des}_{L,R}=\mathbf{e}^{ref}_{L,R}+\triangle\mathbf{e}_{L,R}\in\mathbb{R}^{6}$. The second cost term serves as the desired CAM tracking control. Here, $\mathbf{A}\in\mathbb{R}^{3\times(n+6)}$ denotes the angular component of the centroidal momentum matrix, and $\mathbf{h}^{des}\in\mathbb{R}^{3}$ is the desired CAM calculated from the CP–MPC. The third cost term regulates the optimization variable $\mathbf{\dot{q}}_{1}$. Each term is weighted by its respective positive diagonal weighting matrix: ${\mathbf{w}_{\mathbf{J}}}$, ${\mathbf{w}_{\mathbf{A}}}$, and ${\mathbf{w}_{\mathbf{\dot{q}}}}$. The constraints in (22) consist of two inequality conditions, which limit the joint velocity $\mathbf{\dot{q}}_{1}$ and joint position $\mathbf{q}_{1}$ within their respective boundaries. The vectors $\overline{\mathbf{\dot{q}}}_{1}\in\mathbb{R}^{n+6}$ and $\underline{\mathbf{\dot{q}}}_{1}\in\mathbb{R}^{n+6}$ represent upper and lower bounds for joint velocities, respectively. $\overline{\mathbf{{q}}}_{1}\in\mathbb{R}^{n+6}$ and $\underline{\mathbf{{q}}}_{1}\in\mathbb{R}^{n+6}$ represent upper and lower bounds for joint positions, where $\mathbf{K}_{q}$ is a positive diagonal matrix.
In the second hierarchy, the cost function in (23) consists of two cost terms, similar to those in (21). The first cost term serves as the desired velocity command tracking control, required to maintain the initial pose such as pelvis orientation, chest orientation, and both hand poses. The constraints in (24) include three inequality conditions and one equality condition. The first and second constraints are set to ensure a strict task hierarchy, where lower-priority tasks do not affect higher-priority tasks. If $\mathbf{\bm{\varepsilon}}=0$, CAM control has higher priority than the initial pose recovery since the optimal solution of (23) does not affect the result from the first QP (21). However, it is impossible to generate the return motion without violating the prior CAM control calculated from the higher-level hierarchy. To address this problem, a tolerance $\mathbf{\bm{\varepsilon}}$ was introduced in [23] to relax the strict hierarchy by a specified magnitude. The tolerance $\mathbf{\bm{\varepsilon}}$ is adjusted online according to the 2-norm of the desired CAM. Lastly, the optimized $\mathbf{\dot{q}}$, excluding the underactuated virtual joint, becomes $\mathbf{\dot{q}}^{des}$, while $\mathbf{{q}}^{des}$ is obtained by integrating $\mathbf{\dot{q}}^{des}$ from the current time step.
Note that in our framework, we assigned a higher weighting parameter to maintain the initial pelvis pose compared to the hand and chest pose maintenance, which is a design choice to limit excessive pelvic movement under small disturbances.
@@ VI HQP–based Whole-Body Inverse Kinematics
Figure 4: Simulation results are presented for the robot’s response to external forces applied along both the negative x- and y-directions, shown in (a) and (b) respectively.
VII Results of Simulations and Experiments
VII-A System Overview
@@
This section provides a system overview for both simulations and real robot experiments. The humanoid robot, TOCABI, utilized in the simulations and experiments comprises a total of 33 degrees of freedom (DOFs): 16 for the arms, 12 for the legs, 3 for the waist, and 2 for the neck. Its physical dimensions are approximately 1.8 m in height, weighing around 100 kg, and with a foot size of 15 cm $\times$ 30 cm. The upper-body actuators comprise Parker BLDC motors with harmonic gears, while the lower-body actuators incorporate Kollmorgen BLDC motors with harmonic gears. The current control of the robot is performed by Elmo Motion Control’s Gold Solo Whistle servo controller, and the communication between the servo controller and the main PC is achieved via EtherCAT communication. MicroStrain’s 3DM-GX5-25 IMU is attached to the pelvis, and ATI’s mini85 F/T sensor is mounted on each foot. The algorithmic operation and torque command frequency of the robot are set to 2 kHz, and the walking pattern generator and CP–MPC operate at 50 Hz, due to the high computational load involved, through parallel threads. The walking pattern generator and CP–MPC have respective MPC time horizons of 2.5 s and 1.5 s. It should be noted that the positive direction of the x-axis represents the robot’s forward direction, while the positive direction of the z-axis is the opposite direction of the gravity vector. The simulator employed in this study is the MuJoCo simulator [50]. The qpOASES [51] was employed to solve QP optimization problems, while RBDL [52] was utilized for calculating the kinematics and dynamics of the robot.
Figure 5: Simulation results are presented for the robot’s forward walking while overcoming four unexpected objects.
@@
VII-B Simulations to Validate Robustness of the Proposed Method under External Forces
@@
Simulations were conducted to evaluate the balancing capability of the robot against external forces using the proposed method. The simulations involved a scenario where the robot walked in the place and external forces are applied to the robot’s pelvis link in the x- and y-directions during an SSP (left foot support). The step duration was configured to consist of an SSP of 0.6 s and a DSP of 0.3 s.
@@
In the first simulation, an external force of 240 N was applied to the robot’s pelvis in the negative x-direction for 0.2 s. Fig. 4(a) shows snapshots and data of the robot’s response to the external force. The snapshots illustrate the overall balancing process, where the robot maintains balance through a combination of ZMP control (ankle strategy), CAM control (hip strategy), and back steps (stepping strategy).
@@
The graph in Fig. 4(a) represents the outputs of the CP–MPC and stepping controller in the x-direction. At approximately 5.9 s, the robot was pushed backward, causing the CP error $\xi_{err,x}$ (green line) to increase in the negative x-direction due to the external force. To reduce the CP error, the desired ZMP $z_{x}^{des}$ (red line) is generated by the CP–MPC. Despite the attempts to reduce the CP error through the desired ZMP, it remained high due to the significant external force and the ZMP constraint of -9 cm. As a result, the CP error increased and reached up to -18.3 cm. Then, footstep adjustment $\Delta f_{x,1}$ (blue line) was generated to relax the ZMP constraint and the desired centroidal moment $\tau_{y}^{des}$ (yellow line) was also generated to further reduce the CP error. The desired ZMP was generated up to the relaxed ZMP constraint until approximately 7.6 s, while the centroidal moment was generated up to the limit of -15 Nm until approximately 7.2 s. Meanwhile, the stepping controller performed a total of two back-steps (magenta and cyan line) based on $\Delta f_{x,1}$, moving approximately -23.5 cm in conjunction with an optimal step time $T$ (orange line). As a result, the CP error is reduced to approximately zero, and the robot maintains a balance against the external force.
@@
Fig. 4(b) represents the robot’s data, including the CP–MPC and stepping controllers’ y-direction outputs. At approximately 5.9 s, the external force was applied to the robot in the negative y-direction, leading to a significant deviation in the CP error $\xi_{err,y}$. The CP–MPC attempted to decrease the increasing CP error by generating the desired ZMP $z_{y}^{des}$ up to the ZMP constraint. At approximately 6 s, footstep adjustment of -8.6 cm $\Delta f_{y,1}$ was generated to relax the ZMP constraint, while a desired centroidal moment $\tau_{x}^{des}$ of 15 Nm was also generated to reduce the CP error. Based on the generated $\Delta f_{y,1}$, the stepping controller adjusted the swing foot trajectory and reduced the step time $T$ to allow for stepping at approximately 6.1 s. Subsequently, continuous CP control using each strategy reduced the CP error from approximately -25.7 cm initially, gradually decreasing until the robot was able to maintain balance against external forces.
VII-C Simulation to Validate Robustness of the Proposed Method on Uneven Terrain
In this simulation, the robot walks over objects placed on the ground to validate the balancing performance of the proposed method on uneven terrain. Four objects, each with a length and width of 5 cm, were positioned in front of the robot with a distance of 30 cm between them. To make the robot step on each object alternately, two objects were placed on the left and two on the right with respect to the robot’s center. The thickness of the objects ranged from 1.0 cm to 2.5 cm with an increase of 0.5 cm. The robot walked a distance of 2.0 m with a step length of 0.1 m, and the step duration consisted of 0.6 s of SSP and 0.3 s of DSP.
Fig. 5 shows simulation snapshots and the corresponding output data. In the snapshots, the robot was able to navigate forward while stepping on the objects in its path. The first and second objects were overcome using ZMP and stepping control (ankle and stepping strategies). For the third and fourth objects, CAM control (hip strategy) was additionally employed to navigate past them. The analysis of this process is explained based on the data presented in the graphs. Note that in the second row of Fig. 5, the purple right y-axis highlights the CP error $\xi_{err}$ and footstep adjustment $\Delta\mathbf{f}_{1}$ to enhance the visibility of their variations.
@@ VII-C Simulation to Validate Robustness of the Proposed Method on Uneven Terrain
At the start of the 9th DSP, approximately 9.9 s, the robot encountered the 1.5 cm object. To overcome the resulting disturbance, the CP–MPC generated a footstep adjustment $\Delta f_{y,1}$ of 3.8 cm at approximately 10.8 s, in contrast to the initial encounter with 1.0 cm object. By utilizing $\Delta f_{y,1}$ as a nominal value in the stepping controller, the reference swing foot position $e^{ref}_{L,R,y}$ was adjusted, enabling the robot to overcome the disturbance.
Next, at the start of the 12th DSP, approximately 12.6 s, the robot stepped on the 2 cm object. To alleviate the CP errors, the desired ZMP was generated up to the constraint in both the x- and y-directions. Additionally, $\Delta f_{y,1}$ of -7.0 cm was generated at approximately 13.2 s and $\Delta f_{x,1}$ of -11.2 cm was also generated at 13.5 s. Notably, the large CP error in the y-direction led to a centroidal moment $\tau^{des}_{x}$ of 15 Nm. The robot was able to overcome disturbances by adjusting footstep position and step time to a greater extent than in the second disturbance response, while also implementing CAM control.
Finally, at the start of the 15th DSP, approximately 15.3 s, the robot stepped on the thickest object of 2.5 cm. Due to the large disturbance, the desired ZMP in the x- and y-directions was generated up to each ZMP constraint in three steps. $\Delta f_{y,1}$ of 7.5 cm was generated at approximately 15.9 s, and $\Delta f_{x,1}$ was generated up to -11.4 cm at approximately 16.2 s. The centroidal moment $\tau^{des}_{x}$ in the x-direction was also generated up to the constraint. With footstep position adjustment, step time was also reduced by up to 0.46 s, allowing the robot to overcome the disturbance. In the x-direction, due to the larger range of ZMP control and footstep adjustments compared to the y-direction, only a small amount of centroidal moment $\tau^{des}_{y}$ was generated in overcoming all objects.
In conclusion, this simulation showed that appropriate strategies were executed to mitigate disturbances of various magnitudes caused by uneven terrain.
VII-D Simulation to Evaluate Balancing Performance based on the Different Combinations of Each Balance Strategy
Comparative simulations were conducted to evaluate how the balancing performance varies based on different combinations of balancing strategies in response to disturbances. Presented below is a summary of the implementation methods employed to combine each strategy (see Fig. 6).
  • Method 2) is composed of a combination of ZMP control and stepping control (footstep position and step time adjustment), excluding the CAM control. To achieve this, all terms related to the centroidal moment $\bm{\uptau}$ in (10) and (11) were excluded.
@@
VII-E Simulation to Validate the Effectiveness of the Variable Weighting Method